OpenStudy (anonymous):

The tables below show solutions for two linear equations. If the two equations make up a system of linear equations, in which quadrant is the solution? Quadrant one Quadrant two Quadrant three Quadrant four

OpenStudy (anonymous):

FAN AND MEDAL :)

OpenStudy (anonymous):

If they're linear and you're given several points, you can find the equation of each function by using the formula y=m(x-p)+q.

OpenStudy (anonymous):

You can then create a system of equations by making the two y values equal to one another, I'm guessing.

OpenStudy (anonymous):

The sol'n then should allow you to figure out some values and assign the values a quadrant accordingly. I think that's how you approach this problem.

OpenStudy (anonymous):

??

jimthompson5910 (jim_thompson5910):

are you able to find the equation of each line?

OpenStudy (anonymous):

me? no, I dont really understand what she said...

jimthompson5910 (jim_thompson5910):

how about the slope? are you able to find that?

OpenStudy (anonymous):

um, no... i'm not really good at this sort of thing

OpenStudy (anonymous):

y=mx+b, m is the slope. D:

jimthompson5910 (jim_thompson5910):

are you familiar with the slope formula m = (y2 - y1)/(x2 - x1) ??

OpenStudy (anonymous):

no, i never learned the formulas for this

jimthompson5910 (jim_thompson5910):

for the first table, pick two points and tell me which points you picked

OpenStudy (anonymous):

-7 and 3

jimthompson5910 (jim_thompson5910):

so that's one point (-7,3) what's another?

OpenStudy (anonymous):

-10 and 9

jimthompson5910 (jim_thompson5910):

(-10,9) is another point

jimthompson5910 (jim_thompson5910):

All points are of the form (x,y)

jimthompson5910 (jim_thompson5910):

So the point (-7,3) means x = -7 and y = 3 The point (-10,9) means x =-10 and y = 9

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

hold, on im going to try now

jimthompson5910 (jim_thompson5910):

ok

OpenStudy (anonymous):

i got quadrant 1

jimthompson5910 (jim_thompson5910):

What did you get for the slope?

OpenStudy (anonymous):

linear

jimthompson5910 (jim_thompson5910):

you should have gotten a number for the slope

OpenStudy (whpalmer4):

We have a system of equations. If there's a solution, it's where the lines represented by those equations intersect. Plot the points given, draw a straight line through them. Observe where they intersect.